• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 08:08:13 on October 14,1998

    Entry number 2839

    keyword=shift summary

    Q3 quenched. access to Reset
    Repairs to target completed. EPR/NMR data now look reasonable.
    Began He-3 elastic at 1.73 Gev 9.4M